. If the Ksp for the given reaction is 1.8 x 10 -10  at 250 C, what is the equilibrium concentration of chloride ions, [Cl-]eq, expressed in 10-5 M? 


AgCl(s)      Ag+(aq) + Cl-(aq)

1
Expert's answer
2022-03-12T01:57:17-0500

Solution:

AgCl(s) ⇌ Ag+(aq) + Cl(aq)

__S________S________S____


Thus,

[Ag+] = [Cl] = S


The Ksp expression for AgCl(s) is:

Ksp = [Ag+] × [Cl]

Therefore,

1.8×10–10 = [S] × [S] = S2

[Cl] = S = (1.8×10–10)1/2 = 0.0000134 M = 1.34×10–5 M

[Cl] = 1.34×10–5 M


Answer: The equilibrium concentration of chloride ions (Cl) is 1.34×10–5 M